# Constraining the dispersion measure redshift relation with simulation-based inference

### v1

@@ 1 Introduction
Fast Radio Bursts (FRB) have received significant attention over the past decades, both from cosmological and astrophysical perspectives. First discovered in archival data (Lorimer et al., 2007), these broad, millisecond transient pulses in the radio frequency range get dispersed by free electrons along their line of sight. While their origin is still debated (Petroff et al., 2019) and ranges from Magnetars (Thornton et al., 2013; Bochenek et al., 2020) to binary mergers (Liu et al., 2016), it is clear that the majority of them must be of extragalactic origin due to their highly dispersed signal. The proportionality constant of this dispersion, fittingly called Dispersion Measure (DM), is proportional to the column density of electrons along the line-of-sight.
@@
As a consequence, FRBs have been proposed to be used as a cosmological probe, in particular of the baryon distribution in the Universe. As for all cosmological fields, the electron density can be split into a background and fluctuation component relative to the background. If the host of the FRB is identified, an independent redshift estimate can be obtained. This allows the construction of the DM-$z$ relation, similar to the luminosity distance from supernovae. This relation has been used with current data to measure the baryon density and the Hubble constant (e.g. Zhou et al., 2014; Walters et al., 2018; Hagstotz et al., 2022a; Macquart et al., 2020; Wu et al., 2022; James et al., 2022; Reischke & Hagstotz, 2023a). Likewise, one can study the statistical properties of the DM fluctuations (e.g. Masui & Sigurdson, 2015; Shirasaki et al., 2017; Rafiei-Ravandi et al., 2021; Bhattacharya et al., 2021; Takahashi et al., 2021; Reischke et al., 2021, 2022, 2023).
@@
The ever-increasing number of observed FRBs (currently around 600 unique events, see e.g. Newburgh et al., 2016; CHIME/FRB Collaboration et al., 2021; Khrykin et al., 2024a) leads to raising interest in these events. The Square Kilometre Array (SKA [^1] https://www.skao.int/, Dewdney et al., 2009) should observe $>\!10^{5}$ FRBs. Also, other surveys like DSA-2000 (Hallinan et al., 2019) are planning to detect $>\!10^{4}$ FRBs with host identification. This increasing number makes the modelling and inference process prone to systematic effects. Reischke & Hagstotz (2023b) showed that already with $\sim 300$ FRBs, it becomes necessary to include their covariance to conduct unbiased parameter inference using the DM-$z$ relation, an effect which has been neglected in all studies so far. With around $10^{4}$ FRBs, additional effects such as magnification can become important as well (Takahashi, 2024). Using FRBs as a tool for cosmology and astrophysics therefore requires careful modelling. A lot of these effects can be challenging to model analytically, including the case where systematic effects from the search (or in cosmological terms, survey) strategy will not be tractable.
@@
In this paper, we want to tackle these issues and present simulation-based inference (SBI) of cosmological and astrophysical models via the DM-$z$ relation of FRBs. SBI, sometimes also referred to as Likelihood-Free Inference or Implicit Likelihood Inference, is a Bayesian inference technique that does not require an explicit expression for the likelihood function of the data given the parameters of interest. Instead, the likelihood is implicitly assessed by evaluating the joint probability of the data and parameters from forward simulations that map the parameters to the corresponding synthetic data vectors. This approach offers several advantages over traditional methods that necessitate an explicit form for the likelihood. Firstly, the likelihood can assume any form, thus allowing one to bypass the common assumption of a Gaussian likelihood or the need to define a complex analytical expression for the likelihood. Secondly, for certain models and measurements, it might be impractical or too resource-intensive to determine an analytical likelihood. On the similar side, for SBI, data compression becomes essential for the high-dimensional data and parameter spaces typical in cosmology (Leclercq, 2018; Alsing et al., 2019). The methods available in the SBI framework also vary based on their complexities; from the relatively trivial Approximate Bayesian Computing (ABC, see e.g. Rubin, 1984; Pritchard et al., 1999; Beaumont, 2019) to the latest development in NN. These procedures have been applied to cosmological data analysis (e.g. Fluri et al., 2022; Lu et al., 2023; Lin et al., 2023; Euclid Collaboration et al., 2023; von Wietersheim-Kramsta et al., 2024; Gatti et al., 2024).
@@
The NN requires forward simulations to learn the posterior distribution $p(\boldsymbol{\theta}|\boldsymbol{d})$ where $\boldsymbol{d}$ is simulated given $\boldsymbol{\theta}$. In our case, these forward simulations consist in principle of three components: $(i)$ The large-scale structure (LSS) and the background component are produced using the Generator for Large-Scale Structure (GLASS, Tessore et al., 2023) using halo model power spectra for the three-dimensional electron power spectrum. This will generate log-normal realisations of the electron field with the correct two-point statistics imprinted. $(ii)$ The host contribution, which is simply sampled from a host model probability density function (PDF). $(iii)$ The Milky Way (MW) contribution, for which we will use the standard methods of inferring it from already present electron models (Cordes & Lazio, 2002; Yao et al., 2017; Yamasaki & Totani, 2020). We will then use those forward simulations to train an NN to learn the posterior and sample from the posterior with traditional MCMC. Here, we will use Truncated Sequential Neural Posterior Estimation (TSNPE, Deistler et al., 2022) to conduct the inference within the SBI framework. We aim to fit the amplitude of the DM-$z$ relation and the median and the width of the log-normal host distribution with the available host identified FRBs, providing a roadmap for future cosmological and astrophysical inference with FRBs.
@@ 2.1 FRB basics
The pulses of FRBs undergo dispersion while travelling through the ionized matter distribution in the Universe, leading to a frequency-dependent, $\propto\nu^{-2}$, offset of the bursts’ arrival times. Given this time delay measured as $\delta t(\hat{\boldsymbol{x}},z)$ for an FRB at redshift $z$ in direction $\hat{\boldsymbol{x}}$, the constant of proportionality is the observed dispersion measure: $\delta t(\hat{\boldsymbol{x}},z)=\mathrm{DM}_{\mathrm{tot}}(\hat{\boldsymbol{x}},z)\nu^{-2}$. This DM can be broken up into different components:
@@
The first contribution is $\mathrm{DM}_{\mathrm{LSS}}(\hat{\boldsymbol{x}},z)$, caused by free electrons in the LSS. Here, the dependence on the direction is kept explicitly, since the LSS is correlated. In the literature, $\mathrm{DM}_{\mathrm{LSS}}(\hat{\boldsymbol{x}},z)$ is often split up into an IGM part and a halo part
@@ 2.2 Large-Scale-Structure Contribution
First, we will take a more detailed look at the contribution of the LSS. Quite generally this is given by
@@
For redshifts $z<3$ almost all baryons are ionised, and the DM is therefore rewritten as
@@
where we defined $\mathcal{A}\coloneqq\frac{3c\Omega_{\mathrm{b0}}H_{0}}{8\pi Gm_{\mathrm{p}}}\chi_{\mathrm{e}}\,f_{\mathrm{IGM}}$. The LSS contribution is therefore entirely specified by the statistical properties of the electron density field $\delta_{\mathrm{e}}$. Motivated by numerical simulations of the DM which showed that it follows a log-normal distribution (see e.g. Zhang et al., 2021), we model the electron field using GLASS (Tessore et al., 2023) which can, given a three-dimensional power spectrum of a cosmological field, create log-normal realisations. This is done by dividing the LSS into $N_{\mathrm{shells}}$ non-overlapping and concentric shells. Each shell covers the full sky, spanning the comoving volume between redshift $z_{i}$ and $z_{i+1}$. First, we define a matter weight function along the line-of-sight via:
@@ 2.3 Host Contribution
The next component in the forward simulation is the host contribution, quantifying the effect the host galaxy on the observed DM. Based on the position of the FRB progenitor in the galaxy, the signal may travel through the whole galaxy or parts of it. As such, the induced DM varies accordingly. The effect of a complete or partial travel path through the local host, which translates to high and low DM, is usually described by a log-normal distribution (Macquart et al., 2020; Wu et al., 2022). Other works, e.g. Hagstotz et al. (2022a), have assumed a Gaussian host contribution. Recently, simulations have also shown that a log-normal distribution can fit the host contribution rather well (Theis et al., 2024). With the current data set of FRBs, this choice does not make a difference if priors on the DM are included, as we will show later. For our fiducial case, however, we will choose a log-normal distribution for $\mathrm{DM}_{\mathrm{host,rf}}$ (that is the host contribution in the rest-frame of the host galaxy for which we assume no intrinsic redshift evolution):
@@
where $\mathrm{exp}(\mu)$ and $\mathrm{exp}(2\mu+\sigma_{\mathrm{LN}}^{2})[\mathrm{exp}(\sigma_{\mathrm{LN}}^{2})-1]$ are the median and variance respectively with $x=\mathrm{DM}_{\mathrm{host,rf}}$. The median and the scale ($\sigma_{\mathrm{LN}}$) are free parameters in our study. Note that this samples the rest-frame DM of the host, $\mathrm{DM}_{\mathrm{host,rf}}$.
@@ 2.4 Milky Way contribution
In our analysis, we assume that the MW contribution is accessible from the models described in Cordes & Lazio (2002); Yao et al. (2017); Yamasaki & Totani (2020) and we simply add the numerical value for $\mathrm{DM}_{\mathrm{MW}}(\hat{\boldsymbol{x}})$ in Equation 1. At the current sensitivity level, dictated by the amounts of FRBs available, this addition does not have a sizable influence on the inference of cosmological parameters. However, with more FRBs being observed, it could well be that the addition of the MW contribution leads to a residual correlation in the DM which will be falsely picked up as a cosmological signal. This scenario can in principle be tested with our pipeline, as the MW contribution can simply be added to the simulated DM.
@@ 3.1 Forward simulation
In this section, we describe how we use the previously described ingredients to construct forward simulations for the individual components of Equation 1, which, by simple summation, yield a forward simulation prediction for a set of $\mathrm{DM}_{\mathrm{tot}}(\hat{\boldsymbol{x}}_{a},z_{a})$, with $a=1,\dots,N_{\mathrm{FRB}}$. Here, $\hat{\boldsymbol{x}}_{a}$ and $z_{a}$ are the positions and redshifts of the FRBs from the FRB catalogue. Figure 1 summarises the forward simulation pipeline which goes through the following steps:
@@
This procedure provides a pair ($(\{\theta_{\alpha}\},\{\mathrm{DM}_{\mathrm{tot}}(\hat{\boldsymbol{x}}_{a},z_{a})\})$. Rerunning the pipeline, with parameters sampled from a prior distribution, creates a set of forward simulations, which is used to learn the posterior distribution by the NN. The resolution of the simulation is dependent on the parameter $N_{\mathrm{side}}\in 2^{\mathrm{Z}^{+}}$ as GLASS internally uses HEALPix (Górski et al., 2005). From the flowchart in Figure 1 and the list above, it is clear that any component in the pipeline can easily be exchanged for another model. If, for example, there exists a better model for the host contribution with different parameters, it is easy to replace it.
@@ 3.2 Data compression
The data vector currently is $N_{\mathrm{FRB}}$ dimensional, thus requiring a compression procedure, essentially translating a $d$-dimensional dataset down to $n$ dimensions ($n$ < $d$). This is an essential step, as training the NN with high-dimensional data is slow and can lead to inaccuracies. Lossless data compression preserves the Fisher information of the original data. Therefore, we use the data reduction scheme prescribed in Alsing et al. (2019) reducing the $d$-dimensional data down to a dimension equal to the number of free parameters. Assuming a Gaussian likelihood, this compressed data ($t$) can be obtained as (Tegmark et al., 1997)
@@
where $\boldsymbol{\mu}_{*},\boldsymbol{\boldsymbol{\mathrm{C}}}_{*}$ are the ensemble mean and covariance of the original data at some fiducial value $\boldsymbol{\theta}_{*}$ and $\boldsymbol{\mathrm{d}}$ is the corresponding data vector. $\boldsymbol{\nabla}$ represents the partial derivatives with respect to the free parameters. This fiducial $\boldsymbol{\theta}_{*}$ needs to be optimised to ensure no information loss. Specifically, we use the Fisher scoring method
@@
where $\boldsymbol{{t}}_{k}$ is the compressed statistics for $\boldsymbol{\theta}_{k}$ at the ${k}$-th step. Depending on the complexity of the parameter space, a larger value of $k$ may be required for convergence. In practice, we stop after a finite number of iterations when the increment, the second term on the right-hand side of Equation 29, asymptotically approaches a plateau and further steps are not favoured against the computational time. The components of the Fisher matrix, $\mathrm{F}_{ij}$ are evaluated by its full expression for a Gaussian likelihood:
@@
The three free parameters in our model are the amplitude of the DM-$z$ relation, the median and the scale of the log-normal host distribution. We specify the initial $\boldsymbol{\theta}_{*}$ in Table 1 motivated from Reischke et al. (2022), but we quickly converge to $\boldsymbol{\theta}_{\mathrm{optimal}}=\{0.94,200.74,0.79\}^{\mathrm{T}}$. For the current selection of free parameters, all the derivatives have analytic solutions as the derivative of the covariance of the LSS component (Reischke & Hagstotz, 2023b) scales with $2\mathcal{A}$, with $\mathcal{A}$ being the prefactor in Equation 7. Similarly, the derivatives of the log-normal host covariance are trivial. If we are to fit a more complex model in the future, one would have to calculate those derivatives numerically or for more stable results using a differentiable code. This could for example be achieved by emulating the important quantities.

### v2

@@ 1 Introduction
Fast Radio Bursts (FRB) have received significant attention over the past decades, both from cosmological and astrophysical perspectives. First discovered in archival data (Lorimer et al., 2007a), these broad, millisecond transient pulses in the radio frequency range get dispersed by free electrons along their line of sight. While their origin is still debated (Petroff et al., 2019a) and ranges from Magnetars (Thornton et al., 2013; Bochenek et al., 2020) to binary mergers (Liu et al., 2016), it is clear that the majority of them must be of extragalactic origin due to their highly dispersed signal. The proportionality constant of this dispersion, fittingly called Dispersion Measure (DM), is proportional to the column density of electrons along the line-of-sight.
@@
As a consequence, FRBs have been proposed to be used as a cosmological probe, in particular of the baryon distribution in the Universe. As for all cosmological fields, the electron density can be split into a background and fluctuation component relative to the background. If the host of the FRB is identified, an independent redshift estimate can be obtained. This allows the construction of the DM-$z$ relation, similar to the luminosity distance from supernovae. This relation has been used with current data to measure the baryon density and the Hubble constant (e.g. Zhou et al., 2014; Walters et al., 2018; Hagstotz et al., 2022a; Macquart et al., 2020a; Beniamini et al., 2021; Wu et al., 2022; James et al., 2022b; Reischke & Hagstotz, 2023a). Likewise, one can study the statistical properties of the DM fluctuations (e.g. Masui & Sigurdson, 2015; Shirasaki et al., 2017; Rafiei-Ravandi et al., 2021; Bhattacharya et al., 2021; Takahashi et al., 2021; Reischke et al., 2021, 2022, 2023).
@@
The ever-increasing number of observed FRBs (currently around 600 unique events, see e.g. Newburgh et al., 2016; CHIME/FRB Collaboration et al., 2021; Khrykin et al., 2024a) leads to raising interest in these events. The Square Kilometre Array (SKA [^1] https://www.skao.int/, Dewdney et al., 2009) should observe $>\!10^{5}$ FRBs. Also, other surveys like DSA-2000 (Hallinan et al., 2019) are planning to detect $>\!10^{4}$ FRBs with host identification. This increasing number makes the modelling and inference process prone to systematic effects. Reischke & Hagstotz (2023b) showed that already with $\sim 300$ FRBs, it becomes necessary to include their covariance to conduct unbiased parameter inference using the DM-$z$ relation, an effect which has been neglected in all studies so far. With around $10^{4}$ FRBs, additional effects such as magnification can become important as well (Takahashi, 2024). Using FRBs as a tool for cosmology and astrophysics, therefore, requires careful modelling. A lot of these effects can be challenging to model analytically, including the case where systematic effects from the search (or in cosmological terms, survey) strategy will not be tractable.
@@
In this paper, we want to tackle these issues and present simulation-based inference (SBI) of cosmological and astrophysical models via the DM-$z$ relation of FRBs. SBI, sometimes also referred to as Likelihood-Free Inference or Implicit Likelihood Inference, is a Bayesian inference technique that does not require an explicit expression for the likelihood function of the data given the parameters of interest. Instead, the likelihood is implicitly assessed by evaluating the joint probability of the data and parameters from forward simulations that map the parameters to the corresponding synthetic data vectors. This approach offers several advantages over traditional methods that necessitate an explicit form for the likelihood. Firstly, the likelihood can assume any form, thus allowing one to bypass the common assumption of a Gaussian likelihood or the need to define a complex analytical expression for the likelihood. Secondly, for certain models and measurements, it might be impractical or too resource-intensive to determine an analytical likelihood. On the similar side, for SBI, data compression becomes essential for the high-dimensional data and parameter spaces typical in cosmology (Leclercq, 2018; Alsing et al., 2019). Although not strictly necessary, integrating data compression within the SBI framework can reduce the network training time as the resulting data vector is low-dimensional. Additionally, it reduces the number of simulations required for the inference by effectively reducing the size of the data vector. A downside can be, however, that data compression is not always completely lossless. The inference methods available in the SBI framework also vary based on their complexities, from the relatively trivial Approximate Bayesian Computing (ABC, see e.g. Rubin, 1984; Pritchard et al., 1999; Beaumont, 2019) to the latest development in neural network (NN). Both procedures, SBI and data compression, have been applied to cosmological data analysis with the help of NNs. For example, Lu et al. (2023); Euclid Collaboration et al. (2023) use neural compression to reduce the dimensionality of the data space but follow traditional likelihood analysis on the (NN-estimated) summary statistics, with Fluri et al. (2022) also using ABC on top of this compression. In contrast, Lin et al. (2023); von Wietersheim-Kramsta et al. (2025); Gatti et al. (2024) indeed use SBI with neutral density estimation (NDE) but simpler compression techniques such as score compression.
@@
NDE requires forward simulations to learn the posterior distribution $p(\boldsymbol{\theta}|\boldsymbol{d})$ where $\boldsymbol{d}$ is simulated given $\boldsymbol{\theta}$. In our case, these forward simulations consist in principle of three components: $(i)$ The large-scale structure (LSS) and the background component are produced using the Generator for Large-Scale Structure (GLASS, Tessore et al., 2023) using halo model power spectra for the three-dimensional electron power spectrum. This will generate log-normal realisations of the electron field with the correct two-point statistics imprinted. $(ii)$ The host contribution, which is simply sampled from a host model probability density function (PDF). $(iii)$ The Milky Way (MW) contribution, for which we will use the standard methods of inferring it from already present electron models (Cordes & Lazio, 2002; Yao et al., 2017; Yamasaki & Totani, 2020). We will then use those forward simulations to train a NN to learn the posterior and sample from the posterior with traditional MCMC. Here, we will use Truncated Sequential Neural Posterior Estimation (TSNPE, Deistler et al., 2022) to conduct the inference within the SBI framework. We aim to fit the amplitude of the DM-$z$ relation and the median and the width of the log-normal host distribution with the available host-identified FRBs, providing a roadmap for future cosmological and astrophysical inference with FRBs.
@@ 2.1 FRB basics
The pulses of FRBs undergo dispersion while travelling through the ionized matter distribution in the Universe, leading to a frequency-dependent, $\propto\nu^{-2}$, offset of the bursts’ arrival times (see e.g. Lorimer et al., 2007b; Petroff et al., 2019b). Given this time delay measured as $\delta t(\hat{\boldsymbol{x}},z)$ for an FRB at redshift $z$ in direction $\hat{\boldsymbol{x}}$, the constant of proportionality is the observed dispersion measure: $\delta t(\hat{\boldsymbol{x}},z)=\mathrm{DM}_{\mathrm{tot}}(\hat{\boldsymbol{x}},z)\nu^{-2}$. This DM can be broken up into different components:
@@
The first contribution is $\mathrm{DM}_{\mathrm{LSS}}(\hat{\boldsymbol{x}},z)$, caused by free electrons in the LSS. Here, the dependence on the direction is kept explicitly since the LSS is correlated. In the literature, $\mathrm{DM}_{\mathrm{LSS}}(\hat{\boldsymbol{x}},z)$ is often split up into an IGM part and a halo part
@@ 2.2 Large-Scale-Structure Contribution
First, we will take a more detailed look at the contribution of the LSS. Quite generally, this is given by
@@
For redshifts $z<3$, almost all baryons are ionised, and the DM is, therefore, rewritten as
@@
where we defined $\mathcal{A}\coloneqq\frac{3c\Omega_{\mathrm{b0}}H_{0}}{8\pi Gm_{\mathrm{p}}}\chi_{\mathrm{e}}\,f_{\mathrm{IGM}}$. The LSS contribution is therefore entirely specified by the statistical properties of the electron density field $\delta_{\mathrm{e}}$. Motivated by numerical simulations of the DM which showed that it follows a log-normal distribution (see e.g. Zhang et al., 2021; Walker et al., 2024), we model the electron field using GLASS (Tessore et al., 2023) which can, given a three-dimensional power spectrum of a cosmological field, create log-normal realisations. This is done by dividing the LSS into $N_{\mathrm{shells}}$ non-overlapping and concentric shells. Each shell covers the full sky, spanning the comoving volume between redshift $z_{i}$ and $z_{i+1}$. First, we define a matter weight function along the line-of-sight via:
@@ 2.3 Host Contribution
The next component in the forward simulation is the host contribution, quantifying the effect of the host galaxy on the observed DM. Based on the position of the FRB progenitor in the galaxy, the signal may travel through the whole galaxy or parts of it. As such, the induced DM varies accordingly. The effect of a complete or partial travel path through the local host, which translates to high and low DM, is usually described by a log-normal distribution (Macquart et al., 2020a; Wu et al., 2022). Other works, e.g. Hagstotz et al. (2022a), have assumed a Gaussian host contribution. Recently, simulations have also shown that a log-normal distribution can fit the host contribution rather well (Theis et al., 2024). While the latter study made a quantitative comparison, as the statistical scatter in the simulations is difficult to estimate, a log-normal distribution provides a good fit. Furthermore, with the number of FRBs currently available, this model is accurate enough to capture the complexity of the data. Lastly, the log-normality can also be supported by analytical arguments given, for example, in McQuinn (2014) and Reischke et al. (2024). With the current data set of FRBs, this choice does not make a difference if priors on the DM are included, as we will show later. For our fiducial case, however, we will choose a log-normal distribution for $\mathrm{DM}_{\mathrm{host,rf}}$ (that is the host contribution in the rest-frame of the host galaxy for which we assume no intrinsic redshift evolution):
@@
where $\mathrm{exp}(\mu)$ and $\mathrm{exp}(2\mu+\sigma_{\mathrm{LN}}^{2})[\mathrm{exp}(\sigma_{\mathrm{LN}}^{2})-1]$ are the median and variance respectively with $x=\mathrm{DM}_{\mathrm{host,rf}}$. The median and the scale ($\sigma_{\mathrm{LN}}$) are free parameters in our study. Note that this samples the rest-frame DM of the host, $\mathrm{DM}_{\mathrm{host,rf}}$. Here we assume that the rest-frame host contribution is constant, although there are some hints for redshift evolution (see e.g. Acharya & Beniamini, 2025).
@@ 2.4 Milky Way contribution
In our analysis, we assume that the MW contribution is accessible from the models described in Cordes & Lazio (2002); Yao et al. (2017); Yamasaki & Totani (2020) and we simply add the numerical value for $\mathrm{DM}_{\mathrm{MW}}(\hat{\boldsymbol{x}})$ in Equation 1. At the current sensitivity level, dictated by the amounts of FRBs available, this addition does not have a sizable influence on the inference of cosmological parameters. However, with more FRBs being observed, it could well be that the addition of the MW contribution leads to a residual correlation in the DM, which will be falsely picked up as a cosmological signal. This scenario can, in principle, be tested with our pipeline, as the MW contribution can simply be added to the simulated DM.
@@ 3.1 Forward simulation
In this section, we describe how we use the previously described ingredients to construct forward simulations for the individual components of Equation 1, which, by simple summation, yield a forward simulation prediction for a set of $\mathrm{DM}_{\mathrm{tot}}(\hat{\boldsymbol{x}}_{a},z_{a})$, with $a=1,\dots,N_{\mathrm{FRB}}$. Here, $\hat{\boldsymbol{x}}_{a}$ and $z_{a}$ are the positions and redshifts of the FRBs from the FRB catalogue. Figure 1 summarises the forward simulation pipeline, which goes through the following steps:
@@
This procedure provides a pair ($(\{\theta_{\alpha}\},\{\mathrm{DM}_{\mathrm{tot}}(\hat{\boldsymbol{x}}_{a},z_{a})\})$. Rerunning the pipeline, with parameters sampled from a prior distribution, creates a set of forward simulations, which is used to learn the posterior distribution by the NN. The resolution of the simulation is dependent on the parameter $N_{\mathrm{side}}\in 2^{\mathbb{Z}^{+}}$ as GLASS internally uses HEALPix (Górski et al., 2005). From the flowchart in Figure 1 and the list above, it is clear that any component in the pipeline can easily be exchanged for another model. If, for example, one wants to connect FRB observations closer to cosmological analysis and an effective model of the statistical properties of the gas distribution (e.g. Schneider & Teyssier, 2015; Reischke et al., 2024), this can be easily done.
@@ 3.2 Data compression
The data vector currently is $N_{\mathrm{FRB}}$ dimensional, thus requiring a compression procedure, essentially translating a $d$-dimensional dataset down to $n$ dimensions ($n$ < $d$). This is an essential step, as training the NN with high-dimensional data is slow and can lead to inaccuracies. Lossless data compression preserves the Fisher information of the original data. Therefore, we use the data reduction scheme named score compression prescribed in Alsing et al. (2019), reducing the $d$-dimensional data down to a dimension equal to the number of free parameters. Assuming a Gaussian likelihood, this compressed data ($t$) can be obtained as (Tegmark et al., 1997)
@@
where $\boldsymbol{\mu}_{*},\boldsymbol{\boldsymbol{\mathrm{C}}}_{*}$ are the ensemble mean and covariance of the original data at some fiducial value $\boldsymbol{\theta}_{*}$ and $\boldsymbol{\mathrm{d}}$ is the corresponding data vector. $\boldsymbol{\nabla}$ represents the partial derivatives with respect to the free parameters. We note that the Gaussian likelihood assumption is only employed for the data compression and is removed after that. Hence, the inference itself remains likelihood-free without losing any generality. Improper likelihood assumption can, however, lead to suboptimal compression and this fiducial $\boldsymbol{\theta}_{*}$ needs to be optimised to ensure no information loss (Alsing et al., 2018). Specifically, we use the Fisher scoring method
@@
where $\boldsymbol{{t}}_{k}$ is the compressed statistics for $\boldsymbol{\theta}_{k}$ at the ${k}$-th step. Depending on the complexity of the parameter space, a larger value of $k$ may be required for convergence. In practice, we stop after a finite number of iterations when the increment, the second term on the right-hand side of Equation 29, asymptotically approaches a plateau and further steps are not favoured against the computational time. The components of the Fisher matrix, $\mathrm{F}_{ij}$, are evaluated by its full expression for a Gaussian likelihood:
@@
The three free parameters in our model are the amplitude of the DM-$z$ relation, the median and the scale of the log-normal host distribution. We specify the initial $\boldsymbol{\theta}_{*}$ in Table 1 motivated from Reischke et al. (2022), but we quickly converge to $\boldsymbol{\theta}_{\mathrm{optimal}}=\{0.94,200.74,0.79\}^{\mathrm{T}}$. For the current selection of free parameters, all the derivatives have analytic solutions as the derivative of the covariance of the LSS component (Reischke & Hagstotz, 2023b) scales with $2\mathcal{A}$, with $\mathcal{A}$ being the prefactor in Equation 7. Similarly, the derivatives of the log-normal host covariance are trivial. If we are to fit a more complex model in the future, one would have to calculate those derivatives numerically or, for more stable results, use a differentiable code. This could, for example, be achieved by emulating the important quantities.